OpenStudy (anonymous):

Expand and simplify (i represents the imaginary unit) (3-2i)^7

OpenStudy (anonymous):

I'm sorry, but this question would just take a very long time. Look up Pascal's triangle and use that method to expand this. After, replace the i's with their correct value. Leave i by itself alone, i^2 = -1, i^3 = -i, i^4 = 1.